3. The production function of a firm is described by the following equation 𝑸=𝟏𝟎,πŸŽπŸŽπŸŽπ‘³βˆ’πŸ‘π‘³ 𝟐 where L stands for the units of labor.

a) Draw a graph for this equation. Use the quantity produced in the y-axis and the units of labor in the x-axis.

b) What is the maximum production level?

c) How many units of labor are needed at that point?


1
Expert's answer
2022-03-25T15:07:10-0400

a).


b)maximum production level

ML=MC,where MC =0,then

ML="\frac{dq}{dl}"𝟏𝟎,πŸŽπŸŽπŸŽπ‘³βˆ’πŸ‘π‘³"^2"

=10,000-6l

therefore

10,000-6l=0

6l=10,000

l=1667

Q=10,000l-3l"^2"

=(10000"\times"1667)-(3"\times"1667"^2" )

16,666,666.67-8,336,667

Q=8,329,999.67


c)at maximun,labor

ML=MC,where MC =0,then

ML="\frac{dq}{dl}"𝟏𝟎,πŸŽπŸŽπŸŽπ‘³βˆ’πŸ‘π‘³"^2"

=10,000-6l

therefore

10,000-6l=0

6l=10,000

l=1667